# Terracorder: Sense Long and Prosper

### v1

@@ 2. Prototype and Measurements
We next prototype a Terracorder setup for biodiversity monitoring and validate its long-term sensing capabilities. This application is incredibly important for understanding ecological dynamics; a learning-based scheduler not only enables adaptive scheduling, but also provides useful information for conservationists on large-scale event patterns.
@@
The bare-board active consumption of a Raspberry Pi, as a comparison, can range from $\sim$150mA (Zero) to $\sim$500mA (Model 3); a Raspberry Pi Model 3 setup for acoustic monitoring uses $\sim$300mA on average (sarab, ). PICT, a Pi Zero event-triggered camera setup for biodiversity monitoring, also using a 5MP camera, uses $\sim$150mA in idle operation, and $\sim$374mA with camera/Wifi enabled (pict, ). While the sensors used are not entirely equivalent, their power usage highlights an order of magnitude difference between existing devices and ours.
@@ 3. Isolated Scheduling
We assume a fixed 0.1s recording period on activation to detect ongoing events; if an event is detected, recording continues for the entirety of the event. Otherwise, the device returns to deep-sleep until its next scheduled activation. The simulation also includes an 0.03s delay based on measured Goertzel filtering latency. We assume that all events have equal priority, although this may not be the case in every practical scenario; to address variations, users can extend our reward to quantify the importance of events in each period.
@@
Figure 2. Pareto curve of fixed vs. Q-learning schedules
@@ 4. Collaborative Scheduling
However, learning a distributed scheduling policy across all devices leaves on-device coordination impractical; devices operate independently and are unaware of each other’s states (e.g. overlapping event detections), and the policy is learned over the entire network-space. This necessitates a federated learning approach to inference, and regular data exchange with a base station or external server, which can be resource-intensive and impractical. We instead propose a scheduler that first learns a consistent global policy across the network following eq. 2 (i.e., ignoring detection overlaps and battery-life variability); inference can be performed locally, at minimal overhead, as it only requires observation of the current period $t$ and its detected events. Then, on initial deployment, groups of mutually overlapping devices can establish a round-robin ordering (dist_con, ) rotating every period $t$. The devices can issue an RF ping on every event detection (or per a number of detections). This allows each device to estimate the number of overlapping detections in each of its proximity groups, penalising their activations appropriately based on eq. 3 ($B_{\sigma}$ is not required), unless it’s their slot in the round-robin. This approach not only enables on-device localized fine-tuning and minimizes redundant activations of devices in close proximity but also improves the robustness of the network, as devices learn to adapt their schedules if a neighbor breaks down and stops pinging other devices. The approach can also minimise redundant activations in networks of overlapping devices following non-adaptive schedules.

### v3

@@ 2. Prototype and Measurements
We next prototype a Terracorder setup for biodiversity monitoring and validate its long-term sensing capabilities. This application is incredibly important for building an understanding of ecological dynamics; a learning-based scheduler not only enables adaptive scheduling, but also provides useful information for conservationists on large-scale event patterns.
@@
The bare-board active consumption of a Raspberry Pi, as a comparison, can range from $\sim$150mA (Zero) to ¿500mA (Model 3); a Model 3 setup for acoustic monitoring uses $\sim$300mA on average (sarab, ). PICT, a Pi Zero event-triggered camera setup for biodiversity monitoring, also using a 5MP camera, uses $\sim$150mA in idle operation, and $\sim$374mA with camera/Wifi enabled (pict, ). While the sensors used are not entirely equivalent, their power usage highlights an order of magnitude difference between existing devices and ours.
@@ 3. Isolated Scheduling
We assume a fixed 0.1s recording period on activation to detect ongoing events; if an event is detected, recording continues for the entirety of the event. Otherwise, the device returns to deep-sleep until its next scheduled activation. The simulation also includes an 0.03s delay based on measured Goertzel filtering latency. We assume that all events have equal priority, although this may not be the case in every practical scenario; to address variations, users can extend our reward to quantify event importance in each period.
@@
Figure 2. Performance of fixed vs. Q-learning schedules.
@@ 4. Collaborative Scheduling
However, learning a distributed scheduling policy across all devices leaves on-device coordination impractical; devices operate independently and are unaware of each other’s states (e.g. overlapping event detections), and the policy is learned over the entire network-space. This necessitates a federated learning approach to inference, and regular data exchange with a base station or external server, which can be resource-intensive and impractical. We instead propose a scheduler that first learns a consistent global policy across the network following eq. 2 (i.e., ignoring detection overlaps and battery-life variability); inference can be performed locally, at minimal overhead, as it only requires observation of the current period $t$ and its detected events. Then, on initial deployment, groups of mutually overlapping devices establish clusters over BLE (or another low-power protocol) using GPS or signal strength. The clusters assign a round-robin style ordering between devices, which rotates every period $t$ (dist_con, ). Then, on event detection (or per a number of detections), devices issue a short-range RF ping including a hashed representation of event signal features. This allows each device to estimate the number of overlapping detections in each of its clusters, penalizing their activations appropriately based on eq. 3 ($B_{\sigma}$ is not required), unless it’s their slot in the round-robin. This approach not only enables on-device localized fine-tuning and minimizes redundant activations of devices in close proximity but also improves the robustness of the network, as devices learn to adapt their schedules if a neighbor breaks down and stops pinging other devices. The approach can also minimize redundant activations in networks following non-adaptive schedules.